A particle having angular velocity 10 rad/sec and angular acceleration of 5 rad/sec2 for 10 sec. Find the number of rotations performed by the particle in 10 sec.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From the second law of eqn.
Thetha = ut + 0.5gt^2
thetha = 10*10 + 0.5*5*100 => 100 + 250 => 350*pi/180 radian => 35pi/18.
number of revoultions = 360*18/35pi => 360*18/35*3.14 = 582 revolutions approx.